### v1

# Future-Back Threat Modeling: A Foresight-Driven Security Framework

Abstract
Traditional threat modeling remains reactive—focused on known TTPs and past incident data—while threat prediction and forecasting frameworks are often disconnected from operational or architectural artifacts. This creates a fundamental weakness: the most serious cyber threats often do not arise from what is known, but from what is assumed, overlooked, or not yet conceived, and frequently originate from the future—such as artificial intelligence, information warfare, and supply chain attacks—where adversaries continuously develop new exploits that can bypass defenses built on current knowledge.
To address this mental gap, this paper introduces the theory and methodology of Future-Back Threat Modeling (FBTM). This predictive approach begins with envisioned future threat states and works backward to identify assumptions, gaps, blind spots, and vulnerabilities in the current defense architecture, providing a clearer and more accurate view of impending threats so that we can anticipate their emergence and shape the future we want through actions taken now.
The proposed methodology further aims to reveal known unknowns and unknown unknowns, including tactics, techniques, and procedures that are emerging, anticipated, and plausible. This enhances the predictability of adversary behavior—particularly under future uncertainty—helping security leaders make informed decisions today that shape more resilient security postures for the future.
Keywords Future-Back Threat Modeling $\cdot$ Cyber Threat Foresight $\cdot$ Threatcasting
As first articulated in [isc22025futureback] and later expanded through polycritical foresight discussions [realitystudies2025polycritical], the Future-Back approach integrates epistemic foresight into cyber threat modeling, aligning with ISACA’s revisited perspective on systemic assurance [isaca2025threatmodelingwhitepaper]. Collier (2020) further demonstrated how backcasting can reveal systemic weaknesses within cyber threat intelligence ecosystems, supporting its epistemic role in foresight-driven security [collier2020backcasting].
1 Introduction – The Limits of Past-Forward Security
1.1 Problem Context
Recent industry threat-intelligence reports from CrowdStrike, ENISA, and Mandiant show that adversaries are continuously updating their tactics, techniques, and procedures (TTPs) in ways that emphasise stealth and defence evasion across multiple intrusion vectors [crowdstrike2025gtr, enisa2025tl, mandiant2024mtrends]. CrowdStrike’s Global Threat Report 2025 notes that 2024 saw the emergence of threat actors that “exclusively target cloud environments with unique, cloud-specific skill sets” and that adversaries “strengthened their emphasis on defense evasion in cloud environments, adopting stealth-oriented tactics and tools for initial access and credential access[crowdstrike2025gtr].
ENISA’s Threat Landscape 2025 similarly reports that “phishing continued to be the primary method for initial intrusion” and highlights newly observed ClickFix-style scams that “appeared during the reporting period,” as well as the ClearFake campaign, described as “another innovative technique” that weaponises compromised WordPress sites to distribute infostealers via fake CAPTCHA and verification prompts [enisa2025tl].
Mandiant’s M-Trends 2024 documents how, after Microsoft blocked Office macros by default, attackers adopted “multiple types of phishing payloads and techniques that fell outside of historical norms,” including “code obfuscation, remote payload hosting, placing dropper scripts within archive files, and bypassing email filtering controls,” which Mandiant categorises under “Malware Delivery: Old, Borrowed, and New Techniques[mandiant2024mtrends].
Taken together, these independent observations provide converging evidence that new adversarial TTPs are emerging annually as threat actors adapt to, and often outpace, improvements in organisational defences.
While these reports highlight how adversaries continuously evolve their TTPs, they also underscore a parallel challenge: existing defensive methodologies—particularly threat-modeling frameworks—have not kept pace.
Threat modeling has advanced as a structured way to identify attacker routes and defense mechanisms. Still, most frameworks—STRIDE, PASTA, NIST CSF, ISO—remain limited: they focus on threats visible today, not those emerging from the futures that organizations are actively shaping. This backward-looking approach restricts preparedness against discontinuous risks such as AI-mediated fraud, autonomous exploitation, or geopolitical supply-chain disruptions.
In practice, cybersecurity governance swings between reactive patching and static control checks, with little capacity to predict or prepare for futures beyond the immediate horizon.
Table 1: Comparative Mapping of FBTM Principles Against Major Frameworks and Techniques
Future-Back Threat Modeling Principles NIST CSF 2.0 COBIT ISO 27001:2022 MITRE (ATT&CK / Engage) Honeypots
Using backcasting foresight for defense against unknown unknowns Not explicitly covered (Unique) Not explicitly covered (Unique) Not explicitly covered (Unique) Not explicitly covered (Unique) Not explicitly covered (Unique)
Understanding the adversary and ourselves Identify and detect Governance alignment (risk/control objectives) Identify and detect TTP mapping Intelligence and traps
Use strategic deception as an instrument of epistemic testing Partial:detect anomalies Control assurance and testing Partial:control monitoring Adversary engagement (MITRE Engage) Deception / honeypot traps
Predict the adversary’s next step Threat-informed detection and attack-path analysis Risk scenario analysis Threat intelligence and vulnerability management Tactic chaining / adversary behavior modeling Threat intelligence via honeypot observations
1.2 Theoretical Gap
Strategic foresight and threatcasting methods, especially vanatta2019threatcasting and vescent2018shifting, offer narrative and variable-based tools to anticipate long-term change. However, they usually end before providing actionable engineering artifacts.
On the other hand, backcasting in cyber threat intelligence [collier2020backcasting] aims to reduce hype and clarify perspective, but remains confined to organizational awareness. None of these approaches fully connect future-state insights to the mechanisms of control, telemetry, and design validation that determine whether systems are truly resilient.
1.3 Purpose and Contribution
Future-Back Threat Modeling (FBTM) addresses this gap by integrating foresight directly into the mechanics of threat modeling. Instead of expanding existing threat taxonomies, FBTM re-imagines the epistemology of assurance: every control, policy, or “safety belief” is regarded as a hypothesis that must be falsified under plausible future scenarios.
2 Related Work – Where FBTM Extends Existing Thinking
Traditional threat modeling frameworks like STRIDE, PASTA[pasta_rcTM], NIST CSF guide teams to identify known threats and map controls to reduce risk [microsoft2005stride, nistcsf2024]. However, these frameworks are past-forward: they project known attacker behavior into the future.
In contrast, approaches such as threatcasting [vanatta2019threatcasting] As further conceptual grounding, Vescent and Blakley (2018) explored the role of strategic foresight in evolving security paradigms, providing an early academic basis for integrating foresight into cybersecurity planning [vescent2018shifting].
and strategic foresight [vescent2018shifting] explore possible futures, but stop short of turning those futures into testable security controls.
Similarly, backcasting in cyber intelligence [collier2020backcasting] helps counter hype, yet remains an organizational reflection exercise rather than a technical one.
Future-Back Threat Modeling (FBTM) bridges these worlds: it begins with envisioned futures (foresight), challenges the assumptions that make us feel safe today (epistemic testing), and converts insights into specific technical and governance artefacts (flags $\rightarrow$ gates $\rightarrow$ controls).
In short, FBTM translates imagination into evidence.
3 Foundational Principles of Future-Back Threat Modeling
FBTM rests on five interlocking principles that redefine how organizations think about assurance, foresight, and learning. It moves beyond control-centric threat modeling toward a reflexive discipline that learns by falsifying its own safety beliefs.
3.1 Core Principles of Future-Back Threat Modeling (FBTM)
The Future-Back Threat Modeling (FBTM) framework is grounded in five interrelated epistemic principles that differentiate it from traditional threat modeling. Each principle reframes how security foresight, validation, and governance interact across temporal and cognitive boundaries.
  1. 1.
    Temporal Inversion. FBTM starts from the future and works backward rather than projecting threats forward from the present. It treats a future breach or disruption as if it has already occurred, then reverses the timeline to uncover the fragile assumptions that made it possible. Assumptions—not predefined threats—form the central unit of analysis.
  2. 2.
    Epistemic Stress-Testing. Every control, policy, or belief is viewed as a hypothesis that can be falsified. Security confidence is achieved through disproof rather than confirmation—leveraging deception environments, chaos experiments, and anomaly detection to expose what the system wrongly assumes to be safe.
  3. 3.
    Flags $\rightarrow$ Gates $\rightarrow$ Controls. Signals indicating foresight (flags) trigger governance verification points (gates), which culminate in measurable safeguards (controls). This sequential chain operationalizes foresight, transforming abstract future signals into actionable evidence embedded within architectures, telemetry, and decision workflows.
  4. 4.
    Reflexive Learning and Perspective over Panic. Each disproved assumption enriches the organization’s foresight knowledge base, refining subsequent assumptions and strengthening adaptive resilience. FBTM promotes perspective over panic—ranking futures by systemic impact rather than novelty—to ensure foresight drives proportionate and strategic responses instead of reactive measures.
Figure 1: Conceptual framework of Future-Back Threat Modeling (FBTM). The framework illustrates the cyclic relationship between foresight generation, epistemic testing, organizational learning, and strategic decision inputs. Each step (1–6) represents an iterative layer of hypothesis formulation, validation, and feedback integration.
4 Method Overview – The Future-Back Reflexive Workflow
4.1 Phase I – Foresight and Assumption Mapping
This phase inverts the direction of traditional threat modeling by beginning with future states rather than current systems. The objective is to articulate the beliefs, expectations, and biases that shape an organization’s sense of safety — then make them explicit and testable.
4.1.1 Step 1 – Framing Intelligence Inputs
Gather diverse foresight sources: adversarial simulations, horizon scans, socio-technical trend reports, and geopolitical signals. These are not threat lists, but raw inputs used to construct future scenarios. Output: a set of candidate futures (both desirable and undesirable).
4.1.2 Step 2 – Future-State Assumption Mapping
Within each scenario, identify assumptions that must hold true for the system to remain secure. Classify each as fragile or robust. Each assumption becomes a hypothesis seed — the epistemic unit that drives the next phase.
Table 2: Phase I – Integrated Assumption Mapping (Combining Global and Domain-Specific Foresight)
ID Core Threat (ENISA) Security Assumption (to test) Category Fragility Early Flags / Indicators Business Context & Stakeholder Impact
A1 Supply-Chain Compromise of Software Dependencies "Third-party dependencies and open-source components remain trustworthy; no hidden backdoors infiltrate the build process." Technical / Organizational Fragile Unverified commits from unknown contributors; spike in new or unscanned dependencies; advisory alerts from NVD / GitHub. Global foresight: ENISA warns of systemic OSS compromise. Domain foresight: FinTech platform scaling x10 vendors by 2040; vendor-risk and DevSecOps teams must enforce continuous SBOM validation.
A2 Skill Shortages & Automation Gaps "Operational assurance will remain stable as AI automation replaces human analysts." Organizational / Sociotechnical Fragile SOC backlog >> 2x baseline; false-positive ratio >> 15%; hiring cycle >> 6 months. Global foresight: ENISA flags cyber-talent scarcity. Domain foresight: Company’s SOC headcount frozen while AI tooling expands; HR & CISO accountable for capability resilience.
A3 Artificial Intelligence Abuse / Data Poisoning "AI models and data pipelines remain unpoisoned, auditable, and bias-resilient." Technical / Ethical Fragile Unexplained score drift >> 5%; surge in adversarial prompts; model output bias anomalies. Global foresight: ENISA highlights model manipulation and AI misuse. Domain foresight: AI-driven fraud detection governs >> 80% of transaction approvals by 2040; CDO & AI Board exposure is critical.
A4 Advanced Disinformation & Reputation Attacks "Information-integrity controls can distinguish authentic vs. synthetic narratives affecting the brand." Cognitive / Legal Fragile Deepfake surge >> baseline; coordinated botnet chatter; FIMI pattern detection $\uparrow$. Global foresight: ENISA forecasts political & economic deepfakes. Domain foresight: FinTech brand operating in 30 jurisdictions; Corporate Comms & Legal Risk teams responsible for rapid counter-narratives.
A5 Climate-Driven Infrastructure Disruption "Critical data centres remain operational under regional environmental stress and geopolitical volatility." Operational / Strategic Robust $\rightarrow$ Fragile Regional power instability >> baseline; supply-chain delays >> 2 weeks; disaster-linked downtime >> 4h. Global foresight: ENISA anticipates climate-related infrastructure impact. Domain foresight: Data hubs in Southeast Asia face rising thermal and energy stress; CIO and BC/DR board must ensure adaptive resilience.
This integrated map combines global foresight—EU-level systemic threats identified by ENISA [enisa2023foresight]—with domain-specific foresight derived from the firm’s projected 2040 growth trajectory. Each assumption represents a testable belief about the organization’s future safety conditions. Fragility ratings reveal the intersection between global uncertainty and enterprise-level stressors, indicating where systemic volatility may amplify organizational vulnerabilities.
The top three critical fragile assumptions—A1 (Supply Chain Integrity), A2 (Skill Resilience), and A3 (AI Trustworthiness)—constitute the epistemic core for Phase II, where falsification experiments and governance gates are constructed. These high-fragility nodes mark the transition from conceptual foresight to operational hypothesis testing, anchoring the Future-Back Threat Modeling (FBTM) process in empirically falsifiable ground truth.
5 Phase II – Epistemic Testing and Validation
5.1 Purpose and Rationale
Each safety assumption identified in Phase I is restated as a falsifiable hypothesis, serving as the basis for empirical testing. Instead of confirming that an organization’s controls are effective, FBTM aims to disprove their assumed validity under plausible future conditions. This reversal—testing “what might break our belief in safety” rather than “what sustains it”—creates an epistemic stress test that uncovers hidden vulnerabilities before they develop into actual incidents.
5.2 Step 3 – Hypothesis Framing and Threat Perspective Modeling
Following the epistemic mapping of assumptions in Phase I, this phase introduces a structured, multi-perspective method to reframe each assumption into a testable hypothesis. Drawing from threat intelligence methodologies in NIST SP 800-53 Rev. 5 (2020) and the MITRE ATT&CK framework (2024), threat perspectives are examined through three complementary lenses:
  • Attacker-centric (who): Focusing on adversarial motives, tactics, techniques, and actor profiles.
  • Asset-centric (what/why): Identifying high-value targets and critical data flows within the organization.
  • Architecture-centric (how): Understanding vulnerabilities, interdependencies, and hidden attack paths within systemic design.
This three-dimensional approach goes beyond typical “threat enumeration” to model epistemic exposure—the limits of what the organization believes to be secure. It highlights where safety assumptions have not been tested, especially along high-value operational paths or untested architectural controls.
By aligning these perspectives with future-state assumptions, the framework turns each into a falsifiable statement:
If assumption X is true, system Y stays secure under future Z.
Each hypothesis is recorded in a hypothesis register including:
  1. 1.
    The original assumption;
  2. 2.
    Its related future-state condition;
  3. 3.
    The falsification criteria; and
  4. 4.
    Early signs that may indicate epistemic failure.
This approach is conceptually inspired by ISACA (2025), Threat Modeling Revisited, which suggests a five-stage cycle: goal setting, ecosystem mapping, threat prioritization, mitigation planning, and iterative validation. However, FBTM flips this process: instead of starting from threats and moving to controls, it begins with future assumptions and works backward, using adversarial simulation and hypothesis falsification to identify control weaknesses before designing mitigations. This epistemic inversion sets FBTM apart from traditional forward-looking threat modeling.
5.3 Step 4 – Validation and Experiment Design
The goal is not to confirm existing controls but to identify and challenge the safety assumptions that underpin them. FBTM uses three complementary testing methods to implement this epistemic approach:
  • Deception environments — Using honeypots, decoy credentials, or synthetic data assets as epistemic probes to uncover systemic vulnerabilities and false assurances.
Each test is evaluated with a Foresight-to-Evidence Matrix, which assesses results across three dimensions: technical robustness, organizational responsiveness, and ethical proportionality.
Instead of producing binary results, the analysis yields confidence levels that express probabilistic certainty—how strongly the evidence supports (or refutes) the initial assumption.
This keeps the epistemic focus of FBTM: validation is about systematically falsifying what is presumed to be safe, not proving systems are safe. Therefore, an organization’s foresight strength is measured not by how few assumptions fail, but by how thoroughly it learns from those that do.
6 Phase III – Governance, Executive Perception, and Post-Modeling Action
At this stage, the focus shifts from technical falsification to how those insights reshape governance and decision-making. Future-Back Threat Modeling (FBTM) doesn’t replace frameworks like NIST CSF or RMF—it sits inside them, turning what used to be compliance checklists into learning loops. When a safety assumption fails, the point isn’t blame or proof—it’s translation: how does this new knowledge alter the way we govern, prioritize, and act?
6.1 Step 5a – Governance and Legal Gate
After each epistemic test, results are reviewed through a governance gate that mirrors the Assess–Authorize–Monitor phases of the NIST Risk Management Framework[nist_sp80037r2]. . This isn’t a bureaucratic stop; it’s where discovery becomes accountability.
I usually bring together three lenses in this review:
  1. 1.
    Legal and ethical alignment — ensuring that tests like honeypots or synthetic deception don’t violate data-handling or privacy obligations.
  2. 2.
    Policy connection — mapping what we learned back to existing control frameworks (NIST SP 800-53, ISO 27001, GDPR) [nist2020sp80053, iso27001_2022, gdpr2016].
  3. 3.
    Risk appetite review — asking whether the failed assumption changes what “acceptable” risk looks like at the board level.
When done well, this step turns epistemic evidence into governed risk artifacts: board minutes, updated control objectives, or revised continuous monitoring rules.
6.2 Step 5b – Executive Risk Perception and Decision Economics
In practice, the hardest part of foresight isn’t data—it’s translation. Executives don’t think in control language; they think in trade-offs, timing, and pressure. The goal of this step is to bridge epistemic discovery with executive sense-making.
When future risk becomes visible—say, through a falsified assumption revealed by a honeypot—it enters the same decision triad used in most enterprise frameworks:
  • Mitigate: Adjust defenses, re-prioritize playbooks, or rebalance investment.
  • Accept: Record why this level of uncertainty is tolerable and who owns that decision.
  • Transfer or Avoid: Engage partners, insurers, or modify exposure at a business level.
This aligns directly with NIST CSF “Respond” and “Recover” functions but adds a reflective twist: the conversation isn’t “Is this control compliant?”—it’s “What does this assumption tell us about how we see risk itself?”
6.3 Step 6 – Operational Response and Learning
Once governance decisions are made, the cycle loops back into operations. Here, FBTM connects directly with NIST SP 800-61r2 Incident Response [nist2012sp80061r2]—but treats every incident not just as an event to contain, but as an epistemic experiment to learn from.
When something breaks or behaves unexpectedly:
  • Preparation becomes the documentation of assumptions worth testing.
  • Detection and Analysis focus on evidence of falsification rather than anomaly counts.
  • Containment and Recovery show how fragile design assumptions behave under pressure.
  • Post-Incident Activity transforms those observations into updated foresight scenarios.
Each lesson learned is logged back into the Foresight-to-Evidence Matrix, building a running record of how the organization learns through failure. Over time, these cycles turn what used to be static response plans into a living assurance process—still NIST-aligned, but epistemically aware.
6.4 Closing the Loop
The mark of a mature organization isn’t how few incidents occur, but how quickly it can learn from the ones that do. In FBTM, this loop—Assumption → Hypothesis → Validation → Governance → Response—keeps foresight alive in day-to-day operations. It’s not about predicting the future perfectly; it’s about learning faster than uncertainty unfolds.
7 Phase IV – Synthesis and Strategic Implications
The closing phase of Future-Back Threat Modeling (FBTM) moves from analysis to synthesis—from epistemic testing to strategic sense-making. While the earlier phases examined how assumptions fail, this stage asks how those lessons reshape foresight, policy, and organizational learning.
FBTM’s contribution is not to introduce another risk framework, but to re-architect the flow of knowledge between foresight and control. Instead of treating security as a compliance perimeter, it reframes it as an adaptive system of inquiry. Each epistemic cycle—from assumption to falsification—adds reflexive intelligence to the organization’s risk posture.
Three strategic implications emerge:
  1. 1.
    From Compliance to Cognition: FBTM transforms assurance from a checklist discipline into a cognitive one. Controls become hypotheses to be tested, not rules to be obeyed. This encourages continuous reasoning about what the organization believes to be “safe.”
  2. 2.
    Foresight as a Decision Function: By integrating future-state mapping with governance gates, FBTM embeds foresight directly into executive decision cycles. The focus shifts from predicting discrete threats to rehearsing the organization’s response to uncertainty.
  3. 3.
    Learning as Assurance: In the long term, epistemic learning—the ability to unlearn false confidence—becomes a measure of resilience. The most adaptive organizations are not those with perfect protection, but those that learn the fastest from violated assumptions.
In this sense, FBTM does not stand outside of existing standards like NIST CSF or ISO 27001 [nistcsf2024, iso27001_2022]. Rather, it acts as an epistemic engine inside them, converting static controls into dynamic knowledge.
8 Discussion and Limitations
Like all foresight-driven frameworks, FBTM depends on the quality of human judgment and the contextual richness of its inputs. While it systematizes how organizations learn from failed assumptions, it cannot guarantee that those assumptions are selected or tested without bias. Moreover, the framework’s epistemic focus—valuing falsification over verification—may challenge conventional assurance cultures that prefer certainty and documentation over doubt and discovery.
A second limitation lies in scalability. Applying epistemic stress-testing across a large enterprise demands tooling, telemetry, and cross-domain coordination. Automating this process—for example, through AI-augmented horizon scanning or dynamic assumption registries—remains an open area for exploration.
Finally, empirical validation will require longitudinal studies. Pilot implementations in critical infrastructure or financial sectors could reveal how epistemic learning curves differ from traditional risk maturity models.
Despite these constraints, FBTM provides a conceptual bridge between strategy and security practice—a way to make foresight operational without diluting its intellectual rigor.
9 Conclusion
Future-Back Threat Modeling (FBTM) reframes cybersecurity assurance as a discipline of learning under uncertainty. By moving from foresight to falsification, and from compliance to cognition, it enables organizations to govern risk as a living process rather than a static state.
In doing so, FBTM turns traditional threat modeling on its head: instead of asking “what could attack us,” it asks “what do we wrongly believe to be safe?” Each failed assumption becomes an epistemic success—a new data point in how the organization perceives and adapts to future conditions.
Ultimately, the strength of an organization’s security lies not in how few incidents it suffers, but in how well it learns from those it could not foresee. FBTM provides a language, a method, and a mindset to make that learning deliberate.
Acknowledgments. The author thanks the ISC2 Insights editorial team and foresight practitioners for their early dialogue on epistemic threat modeling.
10 Case Study
Applying the Future-Back Threat Modeling Framework to CVE-2025-64446 Using a Simulated FortiWeb Control-Plane Honeypot.
Table 3: System Components Used in the Future-Back Threat Modeling Case Study for CVE-2025-64446
Component Hardware OS Software / Services Purpose
Honeypot 2 vCPU/2 GB Ubuntu 22.04 DShield, mock API, Python, stunnel4 WAF exposure
API Layer Same host Ubuntu 22.04 Flask API, decoys, logging API emulation
Telemetry Same host Ubuntu 22.04 DShield logs, journal, logger Traffic capture
Network Public IPv4 N/A Port 8443, stunnel4 Attack surface
Figure 2: Simplified honeypot research model used for the CVE-2025-64446 case study.
A lightweight DShield Honeypot was deployed on an Ubuntu server (2 vCPU, 2 GB RAM) and extended to host a simulated FortiWeb management API.
The experimental environment consists of the following components:
  • Attack Surface: A simulated WAF reachable on port 8443.
  • Fake FortiWeb API Endpoints: Non-functional management-plane paths that accept inbound HTTP(S) requests.
  • Honeypot Elements: Decoys and canary fields used to analyze attacker interaction patterns.
  • Logging and Telemetry: Full packet inspection and application-level request capture.
10.1 Real-World Telemetry (Adversary Interaction Evidence)
After exposing the simulated FortiWeb management API, the honeypot received unsolicited external traffic consistent with active exploitation attempts of CVE-2025-64446, which has been reported as exploited in the wild [rapid7_fortiweb_2025].
10.2 Correspondence With Public Reports
The telemetry captured by the honeypot closely matched indicators independently reported by SANS ISC honeypots [sans_fortiweb_2025]. Observed requests included:
  • Access attempts to:
    GET /api/v2.0/cmdb/system/admin/../../../../../cgi-bin/fwbcgi
  • Path-traversal sequences identical to those documented in public reports
  • User-Agent strings commonly used by automated scanners (Firefox 134.x, Safari/537.36, Python-requests)
  • Repeated probing from external IP ranges, including 113.x.x.x
Highlight. Example of unsolicited exploit attempts captured by the deployed honeypot. These requests targeted the simulated FortiWeb management API, matching the path traversal and fwbcgi patterns reported by SANS ISC. The honeypot setup corresponds to the architecture illustrated in Figure 2
Figure 3: Web Honeypot Logs for 2025-11-19 showing probing of ‘/cgi-bin/fwbcgi‘.
10.3 Honeypot-Specific Observations
Within the simulated environment, the honeypot recorded:
  • Full request lines containing the expected fwbcgi exploitation patterns
  • Multiple traversal injections (e.g., ../../../../../)
  • Enumeration attempts across all simulated API endpoints
  • Probes generated from both Linux and Windows-based HTTP clients
The close correspondence between these observations and the SANS dataset provides empirical support for the future-back model’s assumptions regarding attacker behavior toward exposed WAF control-plane surfaces.
The close match between these observations and the SANS dataset provides credible empirical support for the assumptions tested in the future-back model.
10.4 Summary of Framework Integration
The honeypot experiment completes the full Future-Back Threat Modeling (FBTM) cycle. Building on public threat intelligence, a future-state assumption was defined around the compromise of a WAF control-plane interface (Step 2). Attacker-, asset-, and architecture-centric hypotheses were derived from this assumption (Step 3) and tested using a simulated FortiWeb management API deployed within a controlled honeypot environment (Step 4).
Telemetry collected from unsolicited external traffic, including request patterns matching SANS ISC observations of CVE-2025-64446, provided empirical evidence to evaluate and refine these hypotheses. This evidence supports governance adjustments related to the management and exposure of security-control interfaces (Step 5) and informs revisions to the organization’s underlying threat assumptions (Step 6).
In summary, the case study demonstrates how FBTM integrates foresight, hypothesis formulation, epistemic testing, and organizational learning into a coherent, iterative process for improving strategic security posture.